OpenStudy (teller):

Someone mind checking my work? The radioactive substance cesium-137 has a half-life of 30 years. The amount A(t) (in grams) of a sample of cesium-137 remaining after t years is given by the following exponential function. A(t)=381(1/2)^t/30 Find the initial amount in the sample and the amount remaining after 50 years. Round your answers to the nearest gram as necessary. initial amount = 381 after 50 years = 1270

OpenStudy (anonymous):

What do you mean? I see no work. A radioactive substance would have less grams after 50 years. In other words, radioactive substances decreases in grams every year.

OpenStudy (teller):

I sent my work down below where it says initial amount = 381 after 50 years = 1270

OpenStudy (teller):

I plugged the information into my online calculator and that's what I got after 50 years.

OpenStudy (anonymous):

You could at least say you plugged it in. There was no indicator to what you did except put in the answer.

OpenStudy (anonymous):

Because I dont think you plugged it right.

OpenStudy (teller):

A(50)=381(1/2)^50/30 got me 1270

OpenStudy (anonymous):

\[381 \times (0.5)^{\frac{ 50 }{ 30 }}\]

OpenStudy (teller):

But my thing says to use the formula A(t)=381(1/2)^t/30

OpenStudy (anonymous):

Same thing.... (1/2)=0.5

OpenStudy (teller):

120.00748

OpenStudy (teller):

That's what I got.

OpenStudy (teller):

so rounding it would make it 120?

OpenStudy (anonymous):

Yea me too.

OpenStudy (teller):

Alright thank you

OpenStudy (anonymous):

No problem!

OpenStudy (mathmate):

@teller Remember PEMDAS: A(t)=381(1/2)^t/30 equals A(t)=381[(1/2)^t]/30 which is not correct. You need to write A(t)=381(1/2)^(t/30) or else calculators will give wrong numbers. Calculators go strictly with PEMDAS, doesn't know what you want! xD

OpenStudy (anonymous):

Yes exactly. I tried figuring out what @teller did wrong, but couldnt figure on how to get 1270.